# Singular Excitement Beyond the Horizon of a Rotating Black Hole

### v1

@@ III Results
In this section, we present our results for the transition rate of the detector with different black hole angular momenta, boundary conditions of the field, masses of the black hole, and energy gaps of the detector. We also examine the glitches in the rotating case versus the non-rotating case. In all our results, the initial position of the detector is $r_{0}=100\,r_{+}$, and we have chosen the detector to be switched-on when the it starts to freely fall, that is, $\tau_{0}/\ell=0$. Hence, the switch-off time or the total detection time coincides with the proper time of the detector, $\Delta\tau/\ell=\tau/\ell$. Consequently, the detector is at its initial position $r=r_{0}$ when $\Delta\tau/\ell=0$, and it reaches the (geodesic incompleteness) singularity at $r=0$ when $\Delta\tau/\ell=\pi/2$.
@@ III.1 Transition rate for different angular momenta and boundary conditions
Figure 1: The transition rate ˙$\dot{\mathcal{F}}_{\tau}$ of a UDW detector freely falling into a rotating BTZ black hole, with $r_{0}/r_{+}=100$, $M=10^{-4}$, $E\ell=-5$, and (a) $J/M\ell=0$, (b) $J/M\ell=0.5$, and (c) $J/M\ell=0.999$. We calculated the image sum from $n=-N$ to $n=+N$, where $N=200$ in the exterior of the black hole and $N=1000$ behind the outer horizon. The time for the detector to reach the outer horizon is indicated by the vertical solid line, and the time for the detector to reach the inner horizon is indicated by the vertical dashed line. The right edge of the plot is the time to the singularity.
@@
For the values of $M$, $r_{0}$, and $J$ considered in Fig. 1, the behaviour of the transition rate curve changes over different ranges of $\Delta\tau/\ell$, and can be analyzed in three cases: (i) the detector is switched off when it is far away from the black hole, (ii) the detector is switched off when its radial position is comparable to the outer horizon of the black hole, and (iii) the detector is switched off after it has crossed the inner horizon of the black hole.
For short durations of operation, when the detector is far away from the horizon, the transition rate undergoes slow oscillation as dictated by the complex exponential function in the integrand of (16). This oscillation, when computed for the different boundary conditions, will be out of phase for some values of $\zeta$ compared to the others. For $\Delta\tau/\ell$ corresponding to a detector position near the outer horizon of the black hole, the fluctuations in the transition rate become more dramatic. Glitches begin to appear in this region, and after their occurrence, the transition rate curves fluctuate with greater frequency and amplitude. The increased fluctuation arises from the contributions of the $|n|\geq 1$ terms in the image sum–specifically, when the $n$-th term in the image sum encounters a glitch, its behaviour will change from slow and weak oscillation to a more rapid and appreciable oscillation. A detailed discussion on the role of the $|n|\geq 1$ terms is presented in Fig. A1 and Fig. A2 of the appendix.
Outside the inner horizon of the black hole, the transition rate curves are smooth, except for a finite number of glitches. However, if the detector remains on after crossing the inner horizon, it observes a transition rate that is highly irregular. From the insets in Fig. 1 (b) and (c), it can be seen that, past the inner horizon, the transition rate curves become “jagged” for the rotating case, with many inflection points at extremely close times to one another. As the detector continues and approaches the singularity, we see that the transition rate overall increases. We found, however, that only the transition rate function in the static case diverges at the singularity.
@@
Comparing the static case in Fig. 1 (a) with the rotating cases in Fig. 1 (b) and (c), we see that the transition rate curve is similar in all three cases, except for the addition of new behaviours behind the inner horizon when angular momentum is present. The main effect of increasing angular momentum is to push the transition rate curve towards the left (i.e., away from the singularity). We observe that as the angular momentum increases, the detector will encounter the features of the transition rate curve sooner.
@@ III.2 Transition rate for different black hole masses and angular momenta
Figure 2: The transition rate ˙$\dot{\mathcal{F}}_{\tau}$ of a UDW detector freely falling into a rotating BTZ black hole, with $r_{0}/r_{+}=100$, $E\ell=-5$, $\zeta=0$, and (a) $M=10^{-2}$, (b) $M=10^{-4}$, (c) $M=10^{-5}$, and (d) $M=10^{-6}$. We calculated the image sum from $n=-N$ to $n=+N$, where $N=200$ in the exterior of the black hole and $N=1000$ behind the outer horizon. The time for the detector to reach the outer horizon is indicated by the vertical solid black line, and the times for the detector to reach the inner horizon are indicated by the vertical dashed lines, where the colours of the dashed lines match the colours of the black hole angular momenta. The right edge of the plot is the time to the singularity.
@@
For the values of $M$, $r_{0}$, and $J$ considered in the previous subsection, we identified three regions of behaviour for the transition rate function with respect to the detector operating time. To reiterate, these are (i) the far region, where the detector is in the distant exterior of the black hole, (ii) the near outer horizon region, where the detector crosses or approaches close enough to the outer horizon, and (iii) the inner horizon region, where the detector has crossed the second horizon of the black hole. Region (i) is characterized by smooth, slow oscillation of the transition rate function, region (ii) is characterized by larger, faster oscillations interspersed with glitches, and region (iii) is characterized by an extremely high concentration of inflection points, giving the transition rate curve a jagged appearance.
@@
From Fig. 2, it can be seen that detector transition rate is highly sensitive to black hole mass. In particular, region (ii) expands outward (to the left) as black hole mass decreases. Region (i) shrinks as a result, while region (iii) remains confined behind the inner horizon. As region (ii) grows, the detector encounters glitches further and further away from the horizon. Moreover, by adjusting the values of $M$ and $J$, the first type of glitch can be made to appear arbitrarily close to $\Delta\tau/\ell=0$. Thus, for sufficiently small black hole masses, the detection of glitches in the transition rate may serve as an early warning for horizon crossing as proposed in [14].
@@
In the large mass case ($M=10^{-2}$), region (ii) is virtually non-existent. The smooth oscillation of the transition rate function continues until the inner horizon, beyond which irregularities begin to appear. In the medium mass case ($M=10^{-4}$), region (ii) has a relatively small extent, stretching from $\Delta\tau/\ell\approx 1.5$ to the inner horizon. Thus, we see slow oscillation for most of the duration of operation, followed by a relatively brief period of more pronounced oscillations, leading into region (iii). We then considered a transitional mass, $M=10^{-5}$, for which the features of the pronounced oscillations in region (ii) develop appreciably and grow leftward. Finally, in the small mass case ($M=10^{-6}$), the detector traverses through region (ii) for most of its operating duration, the first glitch being encountered as early as $\Delta\tau/\ell\approx 5\pi/32$.
@@
The magnitude of the transition rate increases more quickly beyond the horizon when the black hole mass is small. Far away from the horizon, ˙$\dot{\mathcal{F}}_{\tau}$ oscillates about the value of approximately 0.5, which is true for all four plots in Fig. 2. In the large mass case, the transition rate remains on the order of 1 up to $\Delta\tau/\ell=\pi/2$. However, in the medium mass case, the magnitude of the transition rate grows to roughly 10 in the interval up to $\Delta\tau/\ell=\pi/2$, while in the small mass case, the magnitude grows over 100. Strikingly, we observe again that the transition rate function does not diverge at $\Delta\tau/\ell=\pi/2$ except when $J/M\ell=0$. We note that the smooth appearance of the transition rate curves behind the inner horizon, for the smaller masses, is purely due to the scaling of the plots. When zoomed in sufficiently, one would observe the same jagged shape of the transition rate curve as in the large mass plot.
@@
Finally, we comment on the effect of angular momentum. As was observed in Fig. 1, increasing the angular momentum of the black hole stretches the transition rate curves away from the singularity. Region (iii) grows outward as the angular momentum pushes the inner horizon towards the outer horizon, and region (ii) also stretches outward to some extent. Consequently, keeping all other parameters constant, the same features of the transition rate curve are encountered by the detector at an earlier proper time when the angular momentum is larger.
@@ III.3 Transition rate for different detector energy gaps
Figure 3: The transition rate ˙$\dot{\mathcal{F}}_{\tau}$ of a UDW detector freely falling into a rotating BTZ black hole, with $r_{0}/r_{+}=100$, $J/M\ell=0.999$, $\zeta=0$, and (a) $M=10^{-2}$, (b) $M=10^{-4}$, and (c) $M=10^{-6}$. We calculated the image sum from $n=-N$ to $n=+N$, where $N=200$ in the exterior of the black hole and $N=1000$ behind the outer horizon. The time for the detector to reach the outer horizon is indicated by the vertical solid line, and the time for the detector to reach the inner horizon is indicated by the vertical dashed line. The right edge of the plot is the time to the singularity.
@@
We find that the transition rate curves for all of the energy gaps considered respect the observations regarding black hole mass variation in the previous subsection. Namely, the smaller the mass of the black hole, the further outward region (ii) grows, which holds for all detector energy gaps, as expected since the energy gap does not affect the location of the glitches. Additionally, the magnitude of the transition rate grows more quickly near the singularity when the black hole mass is small, which is observed for any energy gap considered.
@@
As the detector approaches the black hole, the behaviour of the transition rate curves becomes significantly different for different energy gaps, especially when the distance between the detector and the black hole’s outer horizon is small. Although the energy gap does not affect the location of the glitches, its magnitude affects the frequency of the oscillations after each glitch. The insight to be gained is that, when constructing a UDW detector, the value of the energy gap plays a large role in defining what one expects to observe as the detector falls into the black hole.
Figure 4: The transition rate ˙$\dot{\mathcal{F}}_{\tau}$ of a UDW detector freely falling into a rotating BTZ black hole, with $r_{0}/r_{+}=100$, $E\ell=-20$, $\zeta=0$, and $M=10^{-6}$. We calculated the image sum from $n=-N$ to $n=+N$, where $N=200$ in the exterior of the black hole and $N=1000$ behind the outer horizon. The time for the detector to reach the outer horizon is indicated by the vertical solid black line, and the times for the detector to reach the inner horizon are indicated by the vertical dashed lines, where the colours of the dashed lines match the colours of the black hole angular momenta. The right edge of the plot is the time to the singularity.
To further connect these results to our observations in the previous subsection, we show, in Fig. 4, the transition rate ˙$\dot{\mathcal{F}}_{\tau}$ of the detector as a function of the detector’s proper time $\Delta\tau/\ell$ for $E\ell=-20$ (as opposed to $E\ell=-5$ in Fig. 2), given a black hole mass of $M=10^{-6}$. We then compare the transition rates for different black hole angular momenta–specifically, $J/M\ell=0$, 0.5, 0.9, 0.99, and 0.999, as before.
We find that, when the detector is near the outer horizon, the property that increasing angular momentum stretches the transition rate curve to the left is preserved for this new energy gap. However, the behaviour in the small mass case ($M=10^{-6}$) is the most complex, due to the far extent of region (ii). Thus, for example, we see significant discrepancy between the transition rate curves in Fig. 4 at $\Delta\tau/\ell\approx 3\pi/8$, which arises from the highly non-trivial interaction between the oscillatory properties of the transition rate and the effects of glitches.
@@ III.4 Glitches in the rotating BTZ problem
Figure 5: The distribution of glitches in the (a) static case compared to the (b) rotating case with $J/M\ell=0.9$, where we have fixed $M=10^{-4}$ and $r_{0}/r_{+}=100$. We plot detector proper time $\Delta\tau/\ell$ on the vertical axis and term number $n$ on the horizontal axis. The four types of glitches are denoted by different markers. The circle and square markers correspond to glitches that arise from singularities in the terms independent of the field’s boundary conditions in (16), whereas the diamond and triangle markers correspond to glitches that arise from the terms that depend on the boundary conditions. We have extended the vertical axis past the proper time to the singularity for purposes of illustration. The lower solid (dashed) horizontal line indicates the time to the outer (inner) horizon. The upper lines are reflections of the time to the horizons with respect to $\Delta\tau/\ell=\pi/2$.
@@
Recall from section II that the four types of glitches arise from setting $s_{i}^{*}(\tau)=0$ or $s_{i}^{*}(\tau)=\tau$, with the explicit proper times of the glitches given in Eqs (20-23). In the non-rotating case, the glitches are symmetric under $n\rightarrow-n$. In the rotating case, two of the four types of glitches become asymmetric under $n\rightarrow-n$, though they remain symmetric to one another by a reflection through the “origin,” which we take to be ($n=0$, $\Delta\tau/\ell=\pi/2)$. The fact that the glitches are not symmetric as $n\rightarrow-n$ in the rotating case is consistent with the fact that ˙$\dot{\mathcal{F}}_{\tau}$ is in general not symmetric as $n\rightarrow-n$. It is important to note that glitches beyond $\Delta\tau/\ell=\pi/2$ are unphysical, as the detector will have already reached the singularity by then, but we include them in Fig. 5 to illustrate the mathematical structure of the non-differentiable points of the transition rate.
@@
In the rotating case, we also observe that three of the four types of glitches are confined behind the inner horizon. Asymptotically, all glitches (and their reflection with respect to $\Delta\tau/\ell=\pi/2$) accumulate at the inner horizon (or its reflection with respect to $\Delta\tau/\ell=\pi/2$) as $n\rightarrow\pm\infty$. Only one type of glitch (which arises from the first square root for $s_{1}^{*}(\tau)=\tau$) is able to escape from the inner horizon for small $|n|$. Thus, in the limit as $J\rightarrow 0$ for the black hole, and the inner horizon goes to the singularity, we are left with only one observable type of glitch. We find that, despite the richer behaviour of glitches in the rotating BTZ case, we are able to retrieve the non-rotating case in a continuous limit.
@@
Finally, we reiterate that the proper times at which the detector encounters the glitches depends on the initial position of the detector as well as the mass and angular momentum of the black hole. As mass decreases or angular momentum increases, the glitches will stretch downwards in the plots in Fig. 5, meaning that they will be encountered by the detector at earlier proper times. However, the overall shapes of the curves formed by the four types of glitches will remain the same.
III.5 Detector transition rate at the singularity

### v2

@@ III Results
In this section, we present our results for the transition rate of the detector with different black hole angular momenta, boundary conditions of the field, masses of the black hole, and energy gaps of the detector. We also examine the glitches in the rotating case versus the nonrotating case. In all our results, the initial position of the detector is $r_{0}=100\,r_{+}$, and we have chosen the detector to be switched on when it starts to freely fall, that is, $\tau_{0}/\ell=0$. Hence, the switch-off time or the total detection time coincides with the proper time of the detector, $\Delta\tau/\ell=\tau/\ell$. Consequently, the detector is at its initial position $r=r_{0}$ when $\Delta\tau/\ell=0$, and it reaches $r=\frac{r_{-}r_{+}}{r_{0}}=10^{-2}r_{-}$, its distance of closest approach to the (geodesic incompleteness) singularity when $\Delta\tau/\ell=\pi/2$.
@@ III.1 Transition rate for different angular momenta and boundary conditions
Figure 1: The transition rate ˙$\dot{\mathcal{F}}_{\tau}$ of a UDW detector freely falling into a rotating BTZ black hole, with $r_{0}/r_{+}=100$, $M=10^{-4}$, $E\ell=-5$, and (a) $J/M\ell=0$, (b) $J/M\ell=0.5$, and (c) $J/M\ell=0.999$. We calculated the image sum from $n=-N$ to $n=+N$, where $N=200$ in the exterior of the black hole and $N=1000$ behind the outer horizon. The time for the detector to reach the outer horizon is indicated by the vertical solid line, and the time for the detector to reach the inner horizon is indicated by the vertical dashed line. The right edge of the plot is the time to the singularity for $J=0$.
@@
For the values of $M$, $r_{0}$, and $J$ considered in Fig. 1, the behavior of the transition rate curve changes over different ranges of $\Delta\tau/\ell$, and can be analyzed in three cases: (i) the detector is switched off when it is far away from the black hole, (ii) the detector is switched off when its radial position is comparable to the outer horizon of the black hole, and (iii) the detector is switched off after it has crossed the inner horizon of the black hole.
For short durations of operation, when the detector is far away from the horizon, the transition rate undergoes slow oscillation as dictated by the complex exponential function in the integrand of (16). This oscillation, when computed for the different boundary conditions, will be out of phase for some values of $\zeta$ compared to the others. For $\Delta\tau/\ell$ corresponding to a detector position near the outer horizon of the black hole, the fluctuations in the transition rate become more dramatic. Glitches begin to appear in this region, and after their occurrence, the transition rate curves fluctuate with greater frequency and amplitude. The increased fluctuation arises from the contributions of the $|n|\geq 1$ terms in the image sum—specifically, when the $n$-th term in the image sum encounters a glitch, its behavior will change from slow and weak oscillation to a more rapid and appreciable oscillation. A detailed discussion on the role of the $|n|\geq 1$ terms is presented in Fig. A1 and Fig. A2 of the Appendix.
Outside the inner horizon of the black hole, the transition rate curves are smooth, except for a finite number of glitches. However, if the detector remains on after crossing the inner horizon, its transition rate becomes highly irregular. From the insets in Fig. 1 (b) and (c), it can be seen that, past the inner horizon, the transition rate curves become “jagged” for the rotating case, with many inflection points at extremely close times to one another. As the detector continues and approaches the singularity, we see that the transition rate increases overall, but remains finite at the closest approach. We found, however, that only the transition rate function in the static case diverges at the singularity.
@@
Comparing the static case in Fig. 1 (a) with the rotating cases in Fig. 1 (b) and (c), we see that the transition rate curve is similar in all three cases, except for the addition of new behaviors behind the inner horizon when angular momentum is present. The main effect of increasing angular momentum is to push the transition rate curve toward the left (i.e., away from the singularity). We observe that as the angular momentum increases, the detector will encounter the features of the transition rate curve sooner.
@@ III.2 Transition rate for different black hole masses and angular momenta
Figure 2: The transition rate ˙$\dot{\mathcal{F}}_{\tau}$ of a UDW detector freely falling into a rotating BTZ black hole, with $r_{0}/r_{+}=100$, $E\ell=-5$, $\zeta=0$, and (a) $M=10^{-2}$, (b) $M=10^{-4}$, (c) $M=10^{-5}$, and (d) $M=10^{-6}$. We calculated the image sum from $n=-N$ to $n=+N$, where $N=200$ in the exterior of the black hole and $N=1000$ behind the outer horizon. The time for the detector to reach the outer horizon is indicated by the vertical solid black line, and the times for the detector to reach the inner horizon are indicated by the vertical dashed lines, where the colors of the dashed lines match the colors of the black hole angular momenta. The right edge of the plot is the time to the singularity for $J=0$.
@@
For the values of $M$, $r_{0}$, and $J$ considered in the previous subsection, we identified three regions of behavior for the transition rate function with respect to the detector operating time. To reiterate, these are (i) the far region, where the detector is in the distant exterior of the black hole, (ii) the near outer horizon region, where the detector crosses or approaches close enough to the outer horizon, and (iii) the inner horizon region, where the detector has crossed the second horizon of the black hole. Region (i) is characterized by the smooth, slow oscillation of the transition rate function, region (ii) is characterized by larger, faster oscillations interspersed with glitches, and region (iii) is characterized by an extremely high concentration of inflection points, giving the transition rate curve a jagged appearance.
@@
From Fig. 2, it can be seen that the detector transition rate is highly sensitive to black hole mass. In particular, region (ii) expands outward (to the left) as the black hole mass decreases. Region (i) shrinks as a result, while region (iii) remains confined behind the inner horizon. As region (ii) grows, the detector encounters glitches further and further away from the horizon. Moreover, by adjusting the values of $M$ and $J$, the first type of glitch can be made to appear arbitrarily close to $\Delta\tau/\ell=0$. Thus, for sufficiently small black hole masses, the detection of glitches in the transition rate may serve as an early warning for horizon crossing as proposed in [14].
@@
In the large mass case ($M=10^{-2}$), region (ii) is virtually nonexistent. The smooth oscillation of the transition rate function continues until the inner horizon, beyond which irregularities begin to appear. In the medium mass case ($M=10^{-4}$), region (ii) has a relatively small extent, stretching from $\Delta\tau/\ell\approx 1.5$ to the inner horizon. Thus, we see slow oscillation for most of the duration of operation, followed by a relatively brief period of more pronounced oscillations, leading into region (iii). We then considered a transitional mass, $M=10^{-5}$, for which the features of the pronounced oscillations in region (ii) develop appreciably and grow leftward. Finally, in the small mass case ($M=10^{-6}$), the detector traverses through region (ii) for most of its operating duration, the first glitch being encountered as early as $\Delta\tau/\ell\approx 5\pi/32$.
@@
The magnitude of the transition rate increases more quickly beyond the horizon when the black hole mass is small. Far away from the horizon, ˙$\dot{\mathcal{F}}_{\tau}$ oscillates about the value of approximately 0.5, which is true for all four plots in Fig. 2. In the large mass case, the transition rate remains on the order of 1 up to $\Delta\tau/\ell=\pi/2$. However, in the medium mass case, the magnitude of the transition rate grows to roughly 10 in the interval up to $\Delta\tau/\ell=\pi/2$, while in the small mass case, the magnitude grows over 100. We observe again that the transition rate function does not diverge at $\Delta\tau/\ell=\pi/2$ except when $J/M\ell=0$, the situation in which the freely falling detector reaches the singularity. We note that the smooth appearance of the transition rate curves behind the inner horizon, for the smaller masses, is purely due to the scaling of the plots. When enlarged sufficiently, one would observe the same jagged shape of the transition rate curve as in the large mass plot.
@@
Finally, we comment on the effect of angular momentum. As was observed in Fig. 1, increasing the angular momentum of the black hole stretches the transition rate curves away from the singularity. Region (iii) grows outward as the angular momentum pushes the inner horizon toward the outer horizon, and region (ii) also stretches outward to some extent. Consequently, keeping all other parameters constant, the same features of the transition rate curve are encountered by the detector at an earlier proper time when the angular momentum is larger.
@@ III.3 Transition rate for different detector energy gaps
Figure 3: The transition rate ˙$\dot{\mathcal{F}}_{\tau}$ of a UDW detector freely falling into a rotating BTZ black hole, with $r_{0}/r_{+}=100$, $J/M\ell=0.999$, $\zeta=0$, and (a) $M=10^{-2}$, (b) $M=10^{-4}$, and (c) $M=10^{-6}$. We calculated the image sum from $n=-N$ to $n=+N$, where $N=200$ in the exterior of the black hole and $N=1000$ behind the outer horizon. The time for the detector to reach the outer horizon is indicated by the vertical solid line, and the time for the detector to reach the inner horizon is indicated by the vertical dashed line. The right edge of the plot is the time to the singularity for $J=0$.
@@
We find that the transition rate curves for all of the energy gaps considered are consistent with the observations regarding black hole mass variation in the previous subsection. Namely, the smaller the mass of the black hole, the further outward region (ii) grows, which holds for all detector energy gaps, as expected since the energy gap does not affect the location of the glitches. Additionally, the magnitude of the transition rate grows more quickly near the singularity when the black hole mass is small, which is observed for any energy gap considered.
@@
As the detector approaches the black hole, the behavior of the transition rate curves becomes significantly different for different energy gaps, especially when the distance between the detector and the black hole’s outer horizon is small. Although the energy gap does not affect the location of the glitches, its magnitude affects the frequency of the oscillations after each glitch. The insight to be gained is that, when constructing a UDW detector, the value of the energy gap plays a large role in defining what one expects to observe as the detector falls into the black hole.
Figure 4: The transition rate ˙$\dot{\mathcal{F}}_{\tau}$ of a UDW detector freely falling into a rotating BTZ black hole, with $r_{0}/r_{+}=100$, $E\ell=-20$, $\zeta=0$, and $M=10^{-6}$. We calculated the image sum from $n=-N$ to $n=+N$, where $N=200$ in the exterior of the black hole and $N=1000$ behind the outer horizon. The time for the detector to reach the outer horizon is indicated by the vertical solid black line, and the times for the detector to reach the inner horizon are indicated by the vertical dashed lines, where the colors of the dashed lines match the colors of the black hole angular momenta. The right edge of the plot is the time to the singularity for $J=0$.
To further connect these results to our observations in the previous subsection, we show, in Fig. 4, the transition rate ˙$\dot{\mathcal{F}}_{\tau}$ of the detector as a function of the detector’s proper time $\Delta\tau/\ell$ for $E\ell=-20$ (as opposed to $E\ell=-5$ in Fig. 2), given a black hole mass of $M=10^{-6}$. We then compare the transition rates for different black hole angular momenta—specifically, $J/M\ell=0$, 0.5, 0.9, 0.99, and 0.999, as before.
We find that, when the detector is near the outer horizon, the property that increasing angular momentum stretches the transition rate curve to the left is preserved for this new energy gap. However, the behavior in the small mass case ($M=10^{-6}$) is the most complex, due to the far extent of region (ii). Thus, for example, we see a significant discrepancy between the transition rate curves in Fig. 4 at $\Delta\tau/\ell\approx 3\pi/8$, which arises from the highly nontrivial interaction between the oscillatory properties of the transition rate and the effects of glitches.
@@ III.4 Glitches in the rotating BTZ problem
Figure 5: The distribution of glitches in the (a) static case compared to the (b) rotating case with $J/M\ell=0.9$, where we have fixed $M=10^{-4}$ and $r_{0}/r_{+}=100$. We plot detector proper time $\Delta\tau/\ell$ on the vertical axis and term number $n$ on the horizontal axis. The four types of glitches are denoted by different markers. The circle and square markers correspond to glitches that arise from singularities in the terms independent of the field’s boundary conditions in (16), whereas the diamond and triangle markers correspond to glitches that arise from the terms that depend on the boundary conditions. In (a), we have extended the vertical axis past the proper time to the singularity for purposes of illustration. The lower solid (dashed) horizontal line indicates the time to the outer (inner) horizon. The upper lines are reflections of the time to the horizons with respect to $\Delta\tau/\ell=\pi/2$.
@@
Recall from section II that the four types of glitches arise from setting $s_{i}^{*}(\tau)=0$ or $s_{i}^{*}(\tau)=\tau$, with the explicit proper times of the glitches given in Eqs (20-23). In the nonrotating case, the glitches are symmetric under $n\rightarrow-n$. In the rotating case, two of the four types of glitches become asymmetric under $n\rightarrow-n$, though they remain symmetric to one another by a reflection through the “origin,” which we take to be ($n=0$, $\Delta\tau/\ell=\pi/2)$. The fact that the glitches are not symmetric as $n\rightarrow-n$ in the rotating case is consistent with the fact that ˙$\dot{\mathcal{F}}_{\tau}$ is in general not symmetric as $n\rightarrow-n$. It is important to note that glitches beyond $\Delta\tau/\ell=\pi/2$ are unphysical for the nonrotating case, as the detector will have already reached the singularity by then, but we include them in Fig. 5 to illustrate the mathematical structure of the nondifferentiable points of the transition rate.
@@
In the rotating case, we also observe that three of the four types of glitches are confined behind the inner horizon. Asymptotically, all glitches accumulate at the inner horizon (or its reflection with respect to $\Delta\tau/\ell=\pi/2$) as $n\rightarrow\pm\infty$. Only one type of glitch (which arises from the first square root for $s_{1}^{*}(\tau)=\tau$) is able to escape from the inner horizon for small $|n|$. Thus, in the limit as $J\rightarrow 0$ for the black hole, and the inner horizon goes to the singularity, we are left with only one observable type of glitch. We find that, despite the richer behavior of glitches in the rotating BTZ case, we are able to retrieve the nonrotating case in a continuous limit.
@@
Finally, we reiterate that the proper times at which the detector encounters the glitches depends on the initial position of the detector as well as the mass and angular momentum of the black hole. As mass decreases or angular momentum increases, the glitches will stretch downward in the plots in Fig. 5, meaning that they will be encountered by the detector at earlier proper times. However, the overall shapes of the curves formed by the four types of glitches will remain the same.
III.5 Detector transition rate approaching the singularity